OpenStudy (anonymous):

what is the slope of the line that passes through the given points (6,-1) and (-3,-1) a. undefined b. 0 c. -2/3 d. -3/2

OpenStudy (tkhunny):

Look at the y-coordinates and answer the question without doing ANY calculations.

OpenStudy (blurbendy):

hint: slope = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

@tkhunny can you help me

OpenStudy (anonymous):

@blurbendy can you help me

OpenStudy (blurbendy):

yes, what's your question

OpenStudy (anonymous):

can you help me with this question

OpenStudy (blurbendy):

to find the slope of a line that passes through two points, use the formula slope = (y2 - y1) / (x2 - x1) y2 = -2 y1 = -1 x2 = -3 x1 = 6

OpenStudy (tkhunny):

Did you look at the y-coordinates? What is special about them? Can you draw any conclusion?

OpenStudy (isaiah.feynman):

@cutegirl the slope is OBVIOUSLY 0. What does that tell you about the line itself?

OpenStudy (anonymous):

i was not sure lol

OpenStudy (tkhunny):

Since two points determine a line, the x-coordinate is not significant. The y-coordinate will always be -1. This is a horizontal line and the slope is zero. Contract this with a vertical line that has NO slope. This is very different from zero slope.